OpenStudy (hba):

Equation of a circle with ends of diameter at (-3,2) and (5,-6) is :

OpenStudy (hba):

@hartnn @waterineyes I have no clue about this one........

hartnn (hartnn):

that is simple.

hartnn (hartnn):

mid-point is centre

OpenStudy (anonymous):

Find the distance between and divide it by 2 to find radius..

hartnn (hartnn):

distance between the two points = diameter =2r

OpenStudy (anonymous):

Do you know how to find midpoint or not??

OpenStudy (hba):

I know nothing lol but i am going to give you a tutorial on Conic section soon promise :D

OpenStudy (hba):

@waterineyes noooooooo

hartnn (hartnn):

u don't know how to find mid-point ? thats news... just take the average of co-ordinates.

OpenStudy (hba):

x+y/2 but x1 or x2

hartnn (hartnn):

for x-coordinate of midpoint, u take average of x-coordinates of endpoints

hartnn (hartnn):

(-3+5)/2

OpenStudy (anonymous):

Sorry net stopped working..

OpenStudy (anonymous):

And I think your mind too.. @hba

OpenStudy (anonymous):

\[Midpoint(x,y) = (\frac{x_1 +x_2}{2}, \frac{y_1+y_2}{2})\]

OpenStudy (hba):

@waterineyes Yeah right lol and [Math processing error ] Anyways @hartnn I remember (x,y)=(x1+x2/2,y1+y2/2)

OpenStudy (hba):

@waterineyes I cannot see latex

hartnn (hartnn):

so whats the midpoint ? centre

OpenStudy (hba):

(x,y)=(1,-2)

hartnn (hartnn):

thats correct centre

hartnn (hartnn):

now find radius

OpenStudy (anonymous):

Just refresh your page then..

OpenStudy (hba):

So i use the old radius formula ?

OpenStudy (hba):

\[r=\sqrt{g^2+f^2-c}\]

OpenStudy (hba):

@waterineyes @hartnn Do you have a better way ?

hartnn (hartnn):

not that, can u find distance between two points ?

OpenStudy (anonymous):

What are you doing man?? Use distance formula..

OpenStudy (hba):

Yeah

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (anonymous):

\[Distance = \sqrt{(x_2-x_1)^2 - (y_2-y_1)^2}\]

OpenStudy (hba):

\[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (hba):

lol

OpenStudy (anonymous):

Written three times but used not atleast once.. ha ha ha..

hartnn (hartnn):

lol :P

OpenStudy (hba):

\[d=8\sqrt{2}\]

OpenStudy (hba):

what next ?

hartnn (hartnn):

radius = d/2

OpenStudy (hba):

@hartnn ?

hartnn (hartnn):

commented

OpenStudy (anonymous):

This you have got is length of Diameter.. You have to find radius so you will half it..

OpenStudy (hba):

r=4root(2)

hartnn (hartnn):

yup... u know what to do from now....

OpenStudy (anonymous):